A proposed project lasts 3 years and has an initial investment of $500,000. The aftertax cash flows are estimated at $120,000 for year 1, $240,000 for year 2, and $240,000 for year 3. The firm has a target debt/equity ratio of 0.6. The firm's cost of equity is 10% and its cost of debt is 8%. The tax rate is 35%. What is the NPV of this project?
Answer #1
Solution:
NPV of this project is      $5,372.28
Working Notes:
For computation of NPV of the project we have to get discount rate at which cash flow the project are discounted to get their present value and at last Net present value.
Since the firm has target debt equity ratio means, in its capital structure both debt & equity will be , so appropriate discount rate or cost of capital for the project will be Weighted average cost of capital (WACC). So we have to get WACC , then at last NPV for the project.
WACC = (E/V x Ke) + (D/V x After tax Kd)
Where
debt equity ratio = 0.60 or Debt/Equity =0.60/1
Value of equity & debt = V = (1 + debt equity ratio) = 1 + 0.60 =1.60
Weight of Debt = D/V = 0.60 / (1 + 0.60) =0.60/1.60
Weight of Equity =E/V = 1 / (1 + 0.60) =1/1.60
cost of debt = (kd) =8%
Ke = Cost of equity = 10%
Tax rate = 35%
WACC = ??
WACC = (E/V)Ke + (D/V) (1 -Tax rate) Kd
WACC = (1/1.60) x 10% + (0.60/1.60) x (1 -35%) x 8%
WACC = 0.08200
WACC = 8.20%
Now NPV of the Project
Notes: NPV of project we get by discounting project cash flow with WACC , NPV which is positive is acceptable to companies as it shows , the project is able to generate return more than WACC
Year Cash flow PVF @8.20% WACC Present value
0 -500,000 1                              -500,000.00
1 120,000 0.924214418                               110,905.73
2 240,000 0.85417229                               205,001.35
3 240,000 0.789438346                               189,465.20
NPV $5,372.28
Hence The project's NPV is          $5,372.28
Notes: PVF is calculated @ 8.20% = 1/(1+0.082)^n     where n is the period for which PVF is calculated.
Please feel free to ask if anything about above solution in comment section of the question.